OpenStudy (imakeitrainmoney):

What is the solution of 3/8x=6 ?

OpenStudy (studygurl14):

Multiply both sides by 8/3

OpenStudy (studygurl14):

No...

OpenStudy (studygurl14):

Make sure to multiply 6 by 8/3 too

OpenStudy (imakeitrainmoney):

24/3

OpenStudy (studygurl14):

yes, which can be simplified into what?

OpenStudy (imakeitrainmoney):

8

OpenStudy (studygurl14):

yep :)

OpenStudy (studygurl14):

wait, no

OpenStudy (imakeitrainmoney):

thats not 1 of the choices tho..... the choices are 16, 18, 24, or 48

OpenStudy (studygurl14):

6 times 8/3 does not equal 24/3

OpenStudy (studygurl14):

It will be double your original answer, so 16

OpenStudy (imakeitrainmoney):

can u explain how?

OpenStudy (studygurl14):

6 times 8/3 = 48/3 = 16

OpenStudy (imakeitrainmoney):

oh ok i just did my multiplication wrong. thx:)

OpenStudy (studygurl14):

yep :)